Show that if x is a real number, then ⌈x⌉ − ⌊x⌋ = 1 if x is




not an integer and ⌈x⌉ − ⌊x⌋ = 0 if x is an integer.

1
Expert's answer
2021-11-29T05:11:57-0500

1) "x" is integer

"\lceil x\rceil -\lfloor x \rfloor =x-x=0".


2) "x" is not integer

Let "\lfloor x \rfloor=n" , then "n<x<n+1" (or "-n-1<-x<-n" )

Let "\lceil x \rceil =m" , then "m-1<x< m"

We have

"(m-1)+(-n-1)< x-x<m-n"

"m-n-2<0<m-n"

"\lceil x\rceil -\lfloor x \rfloor -2<0< \lceil x\rceil -\lfloor x \rfloor"

"0<\lceil x\rceil -\lfloor x \rfloor <2"

Since "\lceil x\rceil -\lfloor x \rfloor" is integer, it folows that "\lceil x\rceil -\lfloor x \rfloor =1"


So, "\lceil x\rceil -\lfloor x \rfloor 
=\begin{cases}
0, \quad x\in \mathbb{Z}
\\
1, \quad x\not\in \mathbb{Z}
\end{cases}"